Suppose you held a diversified portfolio consisting of a $7,500 investment in each of 20 different common stocks. The portfolio's beta is 1.05. Now suppose you decided to sell one of the stocks in your portfolio with a beta of 1.0 for $7,500 and use the proceeds to buy another stock with a beta of 0.60. What would your portfolio's new beta be? Do not round intermediate calculations. Round your answer to two decimal places.
Answer #1

Decrease in beta as a result of net sales=(1-0.6)=0.4

Hence decrease/stock=0.4/20=0.02

Hence new beta=1.05-0.02

=1.03